OpenStudy (mm3591):

Use the substitution method to solve the system of equations. Choose the correct ordered pair. x + y = 11 -x = -y - 9 A. (9, 2) B. (10, 1) C. (8, 3) D. (11, 0)

OpenStudy (mertsj):

If -x=-y-9 then x = y+9 Replace x in the first equation with y+9 and solve.

OpenStudy (mm3591):

@Mertsj I'm still confused

OpenStudy (loser66):

@mm3591 I confused because I don't know why you are still confused.

OpenStudy (mm3591):

@Loser66 I don't know why I'm confused.

OpenStudy (jdoe0001):

@mm3591 anything with -> \(\large x=-y-9\ then\ x = y+9\)?

OpenStudy (jdoe0001):

well.. shoot missing dash... anyhow... does that make sense? notice -> \(\large -x=-y-9\ then\ x = y+9\)

OpenStudy (jdoe0001):

rats...yeah. notice that

OpenStudy (mm3591):

whats the answer?

OpenStudy (jdoe0001):

@mm3591 yes, no, no the foggiest, even forgot my name, maybe, what day is today?

OpenStudy (jdoe0001):

the answer my friend, is blowing in the wind, the answer is blowing in the wind bob dylan

OpenStudy (mm3591):

NOT HELPFUL!

OpenStudy (jdoe0001):

well... what do you think about the aforementioned? -> x=-y-9 then x = y+9 <--- followed that?

OpenStudy (mm3591):

Confused

OpenStudy (jdoe0001):

rats the minus missing,, darn clipboard anyhow -x=-y-9 then x = y+9

OpenStudy (mertsj):

What did I tell you to substitute for x?

OpenStudy (mm3591):

y+9

OpenStudy (mertsj):

What is the first equation in your problem?

OpenStudy (mm3591):

x + y=11 @Mertsj

OpenStudy (mertsj):

Cross out the x and replace it with y+9

OpenStudy (mertsj):



OpenStudy (mertsj):